# Dinomaly: The Less Is More Philosophy in Multi-Class Unsupervised Anomaly Detection

### v2

@@ A.3 Additional Ablation Studies and Experiments
It is noted that most models are pre-trained with the image resolution of $224\times 224$, except that DINOv2 [30] and DINOv2-R [39] have extra a high-resolution training phase with $518\times 518$. However, directly using the pre-trained weights on a different resolution for UAD without fine-tuning like other supervised tasks can cause generalization problems. Therefore, by default, we still keep the feature size of all compared models to $28\times 28$, i.e., the input size is $392\times 392$ for ViT-Base/14 and $448\times 448$ for ViT-Base/16. Additionally, we train Dinomaly with the low-resolution input size of $224\times 224$. The results are presented in Table A4. Generally speaking, CL+MIM combined models outperform MIM and CL models. In addition, MIM-based models do not benefit from higher resolutions but suffer from them, indicating the lack of generalization on a different input size. Methods involving CL can better adapt to a higher resolution as they optimize the global representation of class tokens in pre-training, which is insensitive to input size. As expected, DINOv2 and DINOv2-R pre-trained on larger inputs can better benefit from higher resolution in Dinomaly. Because some methods, i.e., D-iGPT, DINO, and iBOT, produce similar results to DINOv2 in $224\times 224$, we expect that they also have the potential to be as powerful in Dinomaly if they are pre-trained in high-resolution.
@@
Feature Noise. Prior method [3] proposed to perturb the encoder features by Feature Jitter, i.e. adding Gaussian noise with scale to control the noise magnitude. We evaluate the feature jitter strategy in the proposed Dinomaly by placing it at the beginning of Noisy Bottleneck. As shown in Table A6, both Dropout and Feature Jitter can be a good noise injector in Noisy Bottleneck. Meanwhile, Dropout is more robust to the noisy scale hyperparameter, and more elegant without introducing new modules.
@@
No Noise 98.19 99.55 98.51 97.55 63.11 64.39 93.33
@@
A.4 Limitation
@@
As discussed in section A.1, Dinomaly is used for sensory AD that aims to detect regional anomalies in normal backgrounds. It is not suitable for semantic AD. Previous works have shown that methods designed for sensory AD usually fail to be competitive under semantic AD tasks [3; 2]. Conversely, methods designed for semantic AD do not perform well on sensory AD tasks [62; 37]. Future work can be conducted to unify these two tasks, but according to the "no free lunch" theorem, we believe that methods designed for specific anomaly assumption are likely to be more convincing.

### v5

@@ C Additional Ablation and Experiment
It is noted that most models are pre-trained with the image resolution of $224\times 224$, except that DINOv2 [39] and DINOv2-R [7] have extra a high-resolution training phase with $518\times 518$. Directly using the pre-trained weights on a different resolution for UAD without fine-tuning like other supervised tasks can cause generalization problems. Therefore, by default, we still keep the feature size of all compared models to $28\times 28$, i.e., the input size is $392\times 392$ for ViT-Base/14 and $448\times 448$ for ViT-Base/16. Additionally, we train Dinomaly with the low-resolution input size of $224\times 224$.
The results are presented in Table A1. Within Dinomaly, nearly all foundation models can produce SoTA-level results with image-level AUROC higher than 98%. Generally speaking, CL+MIM combined models outperform MIM and CL models. In addition, most foundations do not benefit from a higher resolution for image-level performance but suffer from it, indicating the lack of generalization on a input size different from pre-training; while as expected, DINOv2 and DINOv2-R pre-trained on larger inputs can better benefit from higher resolution in Dinomaly. Because some methods, i.e., D-iGPT, DINO, and iBOT, produce similar results to DINOv2 in $224\times 224$, we expect that they also have the potential to be as powerful in Dinomaly if they are pre-trained in high-resolution. Employing MAE produces the worst results. MAE was also tested as the backbone of ViTAD[65], resulting in undesirable performances (I-AUROC=95.3), which was attributed to the weak semantic expression caused by the pretraining strategy. It is also noted that MAE is bad in other unsupervised tasks such as ImageNet kNN; therefore, MAE is considered to be less effective in tasks without finetuning.
Table A5: Ablations of Dropout rates in Noisy Bottleneck, conducted on MVTec-AD (%). †: default.
Dropout rate Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
0 (noiseless) 99.19 99.55 98.51 97.55 63.11 64.39 93.33
0.1 99.54 99.75 98.90 98.35 69.46 69.19 94.53
0.2 † 99.60 99.78 99.04 98.35 69.29 69.17 94.79
0.3 99.65 99.83 99.16 98.34 68.46 68.81 94.63
0.4 99.64 99.80 99.23 98.22 67.95 68.33 94.57
0.5 99.56 99.81 99.14 98.15 67.43 67.82 94.64
Table A6: Ablations of reconstruction constraint, conduected on MVTec-AD (%). †: default.
Constraints Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
group-to-group (1 group) 99.64 99.80 99.36 98.18 64.79 65.40 93.96
group-to-group (2 groups)† 99.60 99.78 99.04 98.35 69.29 69.17 94.79
Scalability on VisA and Real-IAD. We demonstrate the performance of different ViT sizes on VisA and Real-IAD in Table A3.
@@
No Noise 99.19 99.55 98.51 97.55 63.11 64.39 93.33
Patch Masking p=0.1 99.27 99.60 98.80 97.92 67.15 66.90 94.18
Patch Masking p=0.2 99.17 99.56 98.59 97.75 66.55 66.32 94.11
Patch Masking p=0.3 99.11 99.54 98.37 97.53 65.48 65.96 93.84
Patch Masking p=0.4 99.20 99.59 98.53 97.71 65.58 66.36 94.15
@@
Table A9: Integrating the essense of Noisy Bottleneck (NB) and Loose Loss (LL) on RD4AD, conducted on MVTec-AD (%). †: Reproduction in our framework; ReLU in ResNet decoder is replaced by GELU, StableAdamW optimizer is used.
Method NB LL Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
RD4AD† 97.8 99.1 97.2 96.4 58.0 59.3 91.9
RD4AD 98.4 99.4 97.9 97.2 58.6 60.4 92.9
RD4AD 98.2 99.2 97.5 96.8 60.0 61.1 92.7
RD4AD 98.5 99.4 97.8 97.2 59.6 61.2 93.0
Table A10: Scaling properties of a previous ViT-based method, ViTAD[65] on MVTec-AD. †: their original setting.
Method Pre-Train Backbone Input Size Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
ViTAD† DINO $256^{2}$ 98.3 99.4 97.3 97.7 55.3 58.7 91.4
ViTAD MAE $256^{2}$ 95.3 97.7 95.2 97.4 53.0 56.2 90.6
ViTAD DINOv2 $256^{2}$ 98.7 99.4 98.1 97.6 55.3 59.1 92.7
ViTAD DINOv2-R $256^{2}$ 98.5 99.3 97.8 97.4 54.5 59.2 92.8
ViTAD† DINO $256^{2}$ 98.3 99.4 97.3 97.7 55.3 58.7 91.4
ViTAD DINO $320^{2}$ 98.3 99.2 97.1 97.6 61.3 63.3 92.4
ViTAD DINO $384^{2}$ 97.8 98.9 96.3 97.5 62.5 63.7 92.4
Table A11: Matching previous methods in computation consumption. Dinomaly can be easily scaled by model size and input size.
Method Params MACs MVTec-AD [3] VisA [70]
I-AUROC P-AUROC P-AUPRO I-AUROC P-AUROC P-AUPRO
DiAD [18] 1331M 451.5G 97.2 96.8 90.7 86.8 96.0 75.2
ReContrast [14] 154.2M 67.4G 98.3 97.1 93.2 95.5 98.5 91.9
RD4AD [10] 126.7M 32.1G 94.6 96.1 91.1 92.4 98.1 91.8
ViTAD [65] 39.0M 9.7G 98.3 97.7 91.4 90.5 98.2 85.1
Dinomaly-Base-$392^{2}$ 148M 104.7G 99.6 98.4 94.8 98.7 98.7 94.5
Dinomaly-Base-$280^{2}$ 148M 53.7G 99.6 98.2 93.6 97.8 98.7 92.4
Dinomaly-Small-$392^{2}$ 37.4M 26.2G 99.3 98.1 94.4 97.9 98.6 93.7
Dinomaly-Small-$280^{2}$ 37.4M 14.5G 99.3 98.0 93.4 96.5 98.5 90.9
@@
Table A13: Performance on MPDD and BTAD under multi-class UAD setting (%). †: method designed for MUAD.
Dateset Method Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
MPDD [24] RD4AD [10] 90.3 92.8 90.5 98.3 39.6 40.6 95.2
SimpleNet [34] 90.6 94.1 89.7 97.1 33.6 35.7 90.0
DeSTSeg [67] 92.6 91.8 92.8 90.8 30.6 32.9 78.3
UniAD [60] 80.1 83.2 85.1 95.4 19.0 25.6 83.8
DiAD [18] 85.8 89.2 86.5 91.4 15.3 19.2 66.1
ViTAD [65] 87.4 90.8 87.0 97.8 44.1 46.4 95.3
MambaAD [17] 89.2 93.1 90.3 97.7 33.5 38.6 92.8
Dinomaly (Ours) 97.2 98.4 96.0 99.1 59.5 59.4 96.6
BTAD [38] RD4AD [10] 94.1 96.8 93.8 98.0 57.1 58.0 79.9
SimpleNet [34] 94.0 97.9 93.9 96.2 41.0 43.7 69.6
DeSTSeg [67] 93.5 96.7 93.8 94.8 39.1 38.5 72.9
UniAD [60] 94.5 98.4 94.9 97.4 52.4 55.5 78.9
DiAD [18] 90.2 88.3 92.6 91.9 20.5 27.0 70.3
ViTAD [65] 94.0 97.0 93.7 97.6 58.3 56.5 72.8
MambaAD [17] 92.9 96.2 93.0 97.6 51.2 55.1 77.3
Dinomaly (Ours) 95.4 98.4 95.6 97.8 70.1 68.0 76.5
Uni-Medical [66] RD4AD [10] 76.1 75.3 78.2 96.5 38.3 39.8 86.8
SimpleNet [34] 77.5 77.7 76.7 94.3 34.4 36.0 77.0
DeSTSeg [67] 78.5 77.0 78.2 65.7 41.7 34.0 35.3
UniAD [60] 79.0 76.1 77.1 96.6 39.3 41.1 86.0
DiAD [18] 78.8 77.2 77.7 95.8 34.2 35.5 84.3
ViTAD [65] 81.8 80.7 80.0 97.1 48.3 48.2 86.7
MambaAD [17] 83.9 80.8 81.9 96.8 45.8 47.5 88.2
Dinomaly (Ours) 84.9 84.1 81.0 96.8 51.7 52.1 85.5
Noisy Rates. We conduct ablations on the discarding rate of the Dropouts in MLP bottleneck, as shown in Table A5. Experimental results demonstrate that Dinomaly is robust to different levels of dropout rate.
Reconstruction Constraint. We quantitatively examine different reconstruction schemes presented in Figure 4. As shown in Table A6, group-to-group LC outperforms layer-to-layer supervision. On image-level metrics, 1-group LC with all layers added performs similarly to its 2-group counterpart that separates low-level and high-level layers; however, 1-group LC mixes low-level and high-level features which is harmful for anomaly localization. More ablations on scalability, input size, pre-trained foundations, etc., are presented in Appendix C.
Noise Bottleneck. UniAD [60] proposed to perturb the encoder features by Feature Jitter, i.e. adding Gaussian noise with scale to control the noise magnitude. It is also easy to borrow the masking strategy of MAE [19] to randomly mask patch tokens before the decoder. We evaluate the effectiveness of feature jitter and patch-masking in Dinomaly by placing it at the beginning of Noisy Bottleneck. As shown in Table A8, both Dropout and Feature Jitter can be a good noise injector in Noisy Bottleneck. Meanwhile, Dropout is more robust to the noisy scale hyperparameter, and more elegant without introducing new modules.
Adaptation on CNN Method. Some proposed elements (Linear Attention and Loose Constraint) are closely bounded to modern ViTs. Loose Loss (hard-mining) can be directly applied to previous CNN-based methods, e.g., RD4AD [10]. Noisy Bottleneck can be adapted to RD4AD with minor modifications (apply dropout before MFF layer). We apply these modules to RD4AD to validate the effectiveness of our contributions. The results are shown in Table A9, where these two elements boost the performance of RD4AD to a whole new level that can be compared with prior MUAD SoTAs.
Scaling of Compared Method. As previously discussed in the Experiment section, compared method cannot fully utilize the scaling of pre-trained method, model size, and input size. For example, RD4AD [10] found WideResNet50 better than WideResNet101 as the encoder backbone. ViTAD [65] found ViT-Small better than ViT-Base. Here, we also present the experiments on pre-training method and input size of ViTAD, as shown in Table A10. It is also noted that the paradigm of ViTAD is very similar to RD4AD (replacing CNN by ViT) as well as the starting point of Dinomaly (the first row in the ablation Table 3).
Computation Comparison. The computation costs of Dinomaly variants were previously presented in Table 4 and Table A2. Here, we compare the computation consumption of Dinomaly and prior works. As shown in Table A11, Dinomaly can be easily scaled by model size and input size to match different application scenarios.
D Additional Dataset
To demonstrate the generalization of our method, we conduct experiments on three more popular anomaly detection datasets under MUAD setting, including MPDD and BTAD and Uni-Medical. The MPDD [24] (Metal Parts Defect Detection Dataset ) is a dataset aimed at benchmarking visual defect detection methods in industrial metal parts manufacturing. It consists of more than 1346 images across 6 categories with pixel-precise defect annotation masks. The BTAD [38] ( beanTech Anomaly Detection) dataset is a real-world industrial anomaly dataset. The dataset contains a total of 2830 real-world images of 3 industrial products showcasing body and surface defects. It is noted that the training set of BTAD is noisy because it contains anomalous samples [25]. Uni-Medical [66] is a medical UAD dataset consisting of 2D image slices from 3D CT volumes. It contains 13339 training images and 7013 test images across three objects, i.e., brain CT, liver CT, and retinal OCT. This dataset is not entirely suitable for evaluating 2D anomaly detection methods, as identifying lesions in medical images requires 3D contextual information. The training hyperparameters are the same to MVTec-AD, except the dropout rate for Uni-Medical is increased to 0.4. The performance of Dinomaly and previous SoTAs is presented in Table A13, where our method demonstrates superior results.
E Results Per-Category
F Qualitative Visualization
G Limitation
@@
As discussed in section A, Dinomaly is used for sensory AD that aims to detect regional anomalies in normal backgrounds. It is not suitable for semantic AD. Previous works have shown that methods designed for sensory AD usually fail to be competitive under semantic AD tasks [60, 10]. Conversely, methods designed for semantic AD do not perform well on sensory AD tasks [42, 43]. Future work can be conducted to unify these two tasks, but according to the ”no free lunch” theorem, we believe that methods designed for specific anomaly assumption are likely to be more convincing.